OpenStudy (anonymous):

The function f is defined by f(x)= see below find f(5x) I have the answer I just need some to check it

OpenStudy (anonymous):

\[\frac{ -2+3x }{ 5+x }\]

OpenStudy (anonymous):

\[\frac{5 \left( 3x-2 \right) }{ x+5 }\]

OpenStudy (calculusxy):

You can evaluate the numerator and denominator and then divide.

ganeshie8 (ganeshie8):

uhmm just replace all "x" wid "5x"

ganeshie8 (ganeshie8):

\(\large f(x) = \frac{ -2+3x }{ 5+x }\) \(\large f(5x) = \frac{ -2+3(5x) }{ 5+(5x) }\)

OpenStudy (calculusxy):

Absolutely :) ^

ganeshie8 (ganeshie8):

u cud have also done it in two steps as @calculusxy mentioned earlier :)

OpenStudy (anonymous):

whoa. ive been working on this all day....geesh...\[3-\frac{ 17 }{5\left( x+1 \right) }\]?

ganeshie8 (ganeshie8):

u wana subtract ?

OpenStudy (anonymous):

\[\frac{ 15x-2 }{ 5\left(x+1 \right) }?\]

ganeshie8 (ganeshie8):

\(\large \color{Red}{\checkmark}\)

ganeshie8 (ganeshie8):

good work !!